OpenStudy (anonymous):

Evaluate the exact value: tan(pi/6)

OpenStudy (anonymous):

sqrt{3}/3

OpenStudy (anonymous):

Steps?

OpenStudy (anonymous):

1/6(180) = 30 tan = sin/cos sin = 1/2 cos = sqrt{3}/2 (1/2)/(sqrt{3}/2) = sqrt{3}/3

OpenStudy (anonymous):

find \[\frac{\pi}{6}\] on the unit circle on the last page of the cheat sheet. second coordinate is sine first coordinate is cosine. take \[\frac{\sin(\frac{\pi}{6})}{\cos(\frac{\pi}{6})}\]

OpenStudy (anonymous):

Oh my goodness satellite! Thank you that's an amazing tool

OpenStudy (anonymous):

i wish i had made it myself! i was starting to and then i found this as thought "why bother" just copy it. if you loose it google "paul's notes" and you will find some good cheat sheets and notes as well

OpenStudy (anonymous):

*lose

jimthompson5910 (jim_thompson5910):

\[\large \tan\left(\frac{\pi}{6}\right)\] \[\large \frac{\sin\left(\frac{\pi}{6}\right)}{\cos\left(\frac{\pi}{6}\right)}\] \[\large \frac{\frac{1}{2}}{\frac{\sqrt{3}}{2}}\] \[\large \frac{1}{2}\times\frac{2}{\sqrt{3}}\] \[\large \frac{1}{\sqrt{3}}\] \[\large \frac{\sqrt{3}}{3}\] So \[\large \tan\left(\frac{\pi}{6}\right)=\frac{\sqrt{3}}{3}\]

OpenStudy (anonymous):

Wonderful, thank you very much. Both of you!